OpenStudy (anonymous):

integrate step by step (x/(x^2 + y^2) ) dy

OpenStudy (lgbasallote):

since it's dy then x and x^2 are just constants \[x \int \frac{1}{x^2 + y^2} dy\] isnt that solveable by one of the arc trig functions now?

OpenStudy (anonymous):

I seen the answer to this problem on wolframalpha but have trouble getting there

OpenStudy (anonymous):

i tried trig sub on this, may be i'm subing the wrong identiy in

OpenStudy (lgbasallote):

 \[\tan \theta = \frac yx\] \[x \tan \theta = y\] \[x\sec^2 \theta d\theta = dy\] \[\sec \theta = \frac{\sqrt{x^2 + y^2}}{x}\] \[x \sec \theta = \sqrt{x^2 + y^2}\] \[x^2 \sec^2 \theta = x^2 + y^2\] so your integral becomes \[x \int \frac{x\sec^2 \theta d\theta}{x^2 \sec^2 \theta}\] \[x \int \frac{d \theta }{x}\] \[\implies \int d\theta\] \[\implies \theta\] \[\implies \tan^{-1}\left( \frac yx \right)\] is that what wolfram got?